A certain preparation consists of liquids x, y, and z in the proportion 5:2:1. How many gallons of the preparation can be made from a stock of materials consisting of 25 gallons of x, 20 gallons of y, and 8 gallons of z?
Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

A certain preparation consists of liquids x, y, and z in the proportion 5:2:1. How many gallons of the preparation can be made from a stock of materials consisting of 25 gallons of x, 20 gallons of y, and 8 gallons of z ?
----------------
Let amount of x be 25 gallons
Since amount of x to z is 5 to 1 amount of z is (1/5)x = 5 gallons
Since amount of y to z is 2 to 1 amount of y is 2z = 10 gallons
=====================
Cheers,
Stan H.

Answer by edjones(8007)

We have to find the liquid that PROPORTIONATELY we have the least of.
start with 8 gallons of z
8*5=40 not enough x. We only have 25 gallons.
8*2=16 enough y.
check y.
20/2*5=50 not enough x
So, the amount we can make is dependent on the amount of x that we have.
Since we have 25 gallons we will use:
25/5=x/1
x=5 gallons of z
25/5=x/2
5=x/2
5*2=2x/2
x=10 gallons of y
I hope that helps.
Ed